CIRSI Project Review

15 April, 1997 C D Mackay.

1. Introduction

This note is a summary of the status of the CIRSI project as of mid-April, 1997. Very considerable progress has been made since the meeting of the Sackler Trustees in October, 1996. Broadly, all areas are progressing well. The first science grade detector has been received and the second is at Heathrow awaiting customs clearance. We had problems with the first engineering grade device which we had also received. It turned out that the part was faulty, after we had spent a great deal of time working on it. Rockwell have since sent another engineering grade device which is working well.

In other areas as well progress has been good. Although there has undoubtedly been some slippage in the detector development schedule, in other areas progress has exceeded expectations, and we remain confident that a working system will be available in good time for the first observing run scheduled at present for Q4, 1997.

2. Major Sub-Systems

2.1 Detector controller

The AstroCam 4100 controller has had the sequencer code modified to allow operation of the HAWAII device. It appears to work well. We still have to implement sub-array read-out, but this is not, in fact, needed for our initial observing runs. We also need to implement and check the multi-quadrant switching system.

2.2 The Controller Interface Unit

This works well now. We intend to re-make the printed circuit board for this unit since the layout could be somewhat improved. This will be done in-house.

2.3 Computer Interface

The PCI interface card works well. The device is operated via simple polled transfers. New device drivers that support full DMA access have recently been received by AstroCam for Windows 95. The manufacturers of the interface card have now delivered full Linux drivers (including all source code). They appear to work, and so it is likely that we will use a Linux (Unix on PC) environment for the camera control.

2.4 HAWAII device Operations

The devices are mounted in a modular printed circuit assembly with cryogenic compatible components and connectors. It allows the four devices to be close packed inside the main dewar. It is shown below:

The silicon multiplexer has now worked well for some time. This has allowed all the imaging software we have (PixCel from AstroCam) to be run to check device operation. Images have been successfully obtained from the multiplexer inside the test dewar at liquid nitrogen temperatures (77K). The performance was excellent.

We also tried to operate the original engineering grade device without success. There is evidence of damage that may have occurred at Rockwell. Unfortunately not much information is given about the device characteristics for diagnostic purposes. Rockwell tell me that the device was in fact the second device ever tested, in May 1994. It has kicked around the lab since then, and was not re-tested since then. They have no idea whether it was subjected to some nasty experience. They agreed to send a replacement device which we received in early April, 1997. They confirmed that what we were doing seemed to be correct and that we really should have every success with it.

We have now been able to operate it successfully at cryogenic temperatures. Images have been taken, and one of the first is shown here:

There is still more work to do with this test device but we hope to begin operating the first science grade device within a few weeks, once some changes have been made to the system electronics board.

The first science grade device has been received and substantially exceeds the specification for a full-price, top-grade device. The second science grade device is at Heathrow awaiting customs clearance. It also substantially exceeds the specification for a full-price, top-grade device.

No attempt will be made to operate this device until an engineering grade device has been operated successfully on many occasions.

2.5 Test Dewar

This has been operated cold successfully on several occasions with the multiplexer device. It easily reaches the desired temperature. At present the filter wheel mechanism needs improving. Now that we are able to use an IR sensitive device we can check for light leaks and dark current, but everything looks fine at present. A new filter wheel is under construction now.

2.6 System Testing hardware and software

We have all necessary calibration equipment and software now in place to allow us to press ahead with this as soon as we have a working engineering grade device. We have a 950nm filter for use with a calibrated light source, but still await the H, J and other near IR filters.

2.7 Filters

We have had great difficulties with our filter supplier (mostly bad luck and naiveté on their part) not being able to meet his own specification. Because of worries here we have ordered other sets from another supplier, and expect deliveries in a few weeks. We already have a 950 nm filter for carrying out the initial photometric tests.

2.8 Science Dewar

This has progressed well. The main housing has been completed. The internal liquid nitrogen canister and its supports is under construction, and the internal cooled filter wheel design is complete and awaiting manufacture. We will probably put this into a commercial machine shop because of the overload of the engineering department workshops (which has been largely caused by the Deep Sky Initiative). The detector mounts are already in use in the test dewar, and other internals are well ahead in the design process.

2.9 Acquisition and Guiding System

The camera system has been purchased from AstroCam and works well. The feed optics have been designed, and the pick-off mirror internal to the Science Dewar has been designed. The software to drive the camera in sub-array read-out mode works.

We have done nothing towards a full auto-guider system. This will be done as time permits since manual guiding is likely to be entirely satisfactory for the first runs. For the La Palma telecsopes that we expect to use for our first runs, unguided operation is almost certainly possible.

2.10 Telescope Mounts

This awaits a decision by the time allocation committees as to which telescope(s) we will use. We hope to use the Isaac Newton 2.5 metre Telescope and the William Herschel 4.2 metre Telescope, both on La Palma. Martin Beckett is scheduled to pay a short visit to these telescopes in May to assess mounting and any other potential problems that there might be with operating there.

2.11 Software

All software that we need for running the camera and characterising it as a detector system already exists under Windows 95 on a PC.

The software for running the camera system at the telescope will run on a PC running Linux (Unix). We shall use a TCL/Tk interface to drive the camera control library (41UI) which already runs under Linux. We have used the toolboxes and widget libraries that are available from places such as the European Southern Observatory (complete with documented source code) that provide a substantial interactive user interface under X-Windows on the Linux platform.

A detailed specification of the functionality of that package and a listing of the components needed and/or available is in preparation (by MGB). This will allow the camera to be operated for multi-frame fast read-out, with data saving in FITS format, fast on-line co-addition of identical frames, and the ability to read multiple non-destructive frames (more for spectroscopic operation).

The next software phase requires work on the pre-processing pipeline where the image dither is removed, and flat-fielding and bad pixel removal will occur. This is to be done by RGM and associates.

3. Conclusions

Progress continues at a satisfactory pace. The main issues that we are working to push ahead with relate to the need to recover some of the time lost in getting the HAWAII arrays working properly, and the need to accelerate some of the parts manufacture that are slightly bogged down at the Engineering laboratories.

However, there is still significant contingency in the schedule to ensure we will be able to achieve our fundamental objective to be operational on a telescope in the last quarter of 1997.

Published by Craig D Mackay (, Project Manager, 15 April 1997.